# Object-centric proto-symbolic behavioural reasoning from pixels

### v2

@@ 5.2 Grounded protosymbols and preferences in the latent space
A key feature of OBR is that it learns grounded representations and manipulates them in the latent manifold, as well as generating the agent preferences (goals) in the object-centric latent representation. Agents’ internal goals are tied to the world and the actions that it can perform in it. Hence, bringing benefits but also restrictions as it forces the system to perform embodied reasoning. Learned protosymbolic rules depend on the potential learnable object features and how the agent can act on them—See latent space traversals in Appendix 8. For instance, the agent cannot learn to change the color of an object if all objects are in grey-scale or there is not a policy that can cause the desired effect. As a consequence, conditional behaviours that cannot be interpreted in the objects features space cannot be properly resolved in execution. On the good side, goal-conditioned behaviours are always grounded and therefore the gap between reasoning and control vanishes.
@@ 5.3 Shortcomings and future work
Following the brain inspiration, we employed an iterative inference procedure in order to benefit from postdictive inference (“smoothing”) in a sliding window approach. The resulting architecture is limited to the accuracy of the perception module. Thus, while this work focused on the proof-of-principle of the advantage of object-centric representations in behavioural reasoning, more powerful and sophisticated architectures (e.g. vision transformers) can be used to improve perceptual inference and scale to naturalistic images. Second, in the experiments, the effect of the real actions in the latent representation is assumed to be linear. While this could fit a hierarchical interpretation of cognition, it prevents to properly manipulate complex non-linear dynamics. While OBR can recover from collisions, it cannot plan ahead with them. Further work should investigate complex collisions, for instance, by incorporating other inductive biases, such as interaction networks [16] or sparsely changing latent representations to encode relevant events [48]. Third, the achieved agent behaviours are much less complex than by current SOTA LLMs/VLM approaches such as RT architectures [49]. OBR provides, however, goal-conditioned behavioural reasoning, which is a major challenge in the field [50]. Increased expressivity of the reasoning should be investigated, for instance, by using OBR as the interface between symbolic (e.g., graph-based, language) and subsymbolic representations. Finally, and very relevant to operationalize this approach for robotics applications the agent needs to include the body restrictions of performing an action in the world [51, 29]. This can be done by changing the action field mapping by a robotic arm controller.
@@ Acknowledgments
This work has been funded by the SPIKEFERENCE project, co-funded by the Human Brain Project (HBP) Specific Grant Agreement 3 (ID: 945539).
@@ 6.1.1 Perceptual inference module
This module used a latent dimension of 16. Note that, in the output of the refinement network, this number is doubled once as each latent belief is encoded by a mean and variance, and then doubled again as we represent (and infer) both the states and their first-order derivatives. In the decoder, the latent dimension is doubled only once, as the state derivatives do not enter into the reconstruction of a video frame. As in [10], we use a spatial broadcast decoder, meaning that the latent beliefs are copied along a spatial grid with the same dimensions as a video frame, and each latent vector is concatenated with the $(x,y)$ coordinate of its grid location, before passing through a stack of transposed convolution layers. Decoder and refinement network architectures are summarized in the tables below. The refinement network takes in 16 image-sized inputs, which are identical to those used in [10], except that we omit the leave-one-out likelihoods. Vector-sized inputs join the network after the convolutional stage (which processes only the image-sized inputs), and consist of the variational parameters and (stochastic estimates of) their gradients.
Decoder (states)
Type Size/#Chan. Act. func. Comment Input (${\bm{\mathbf{\lambda}}}$) 32 Broadcast 34 Appends coordinate channels ConvT $5\times 5$ 32 ELU ConvT $5\times 5$ 32 ELU ConvT $5\times 5$ 32 ELU ConvT $5\times 5$ 32 ELU ConvT $5\times 5$ 4 Sigmoid or Softmax Outputs RGB + mask
Refinement network (states)
Type Size/#Chan. Act. func. Comment Linear 64 LSTM 128 tanh Concat $[...,{\bm{\mathbf{\lambda}}},\nabla_{\bm{\mathbf{\lambda}}}\mathcal{L}]$ 256 Appends vector-sized inputs Linear 128 ELU Flatten 800 Conv $5\times 5$ 32 ELU Conv $5\times 5$ 32 ELU Conv $5\times 5$ 32 ELU Inputs 16
@@
To enable action planning across episodes of non-trivial length, it is important that OBR’s perceptual inference and predictive abilities remain stable across longer time windows. In addition, a great benefit of OBR’s slot-based architecture is that object slots can be added or removed at will, without having to learn additional connection weights [36]. However, it is not a given that performance will be robust to such variations. Here, we test the robustness of OBR’s world model, as measured by segmentation and reconstruction accuracy, to both of these dimensions (Fig. 9). Having been trained on 4-frame videos with 3 objects, OBR generalizes well to longer videos with fewer or more objects present in the scene. A slight drop-off in performance towards later frames in the 12-frame testing videos is easily remedied by a very short bout (3 epochs) of additional training with an additional set of longer (also 12-frame) training videos. Generalization to different numbers of objects is characterized by a slight drop in performance as the number increases, but this might also be (partly) attributed to increased complexity of the resulting images.
@@ 6.1.2 Action inference module
A key challenge in multi-object environments such as the evaluated environment, is that an agent’s internal representation of objects has an unknown (and possibly imperfect) correspondence to the objects in the environment. Even if the object features are inferred with perfect accuracy, the order of the objects in the representation is arbitrary. To solve this correspondence problem, we define an action space in which accelerations can be placed at pixel locations within the environment’s image grid, and objects receive the sum of all accelerations that coincide with their visible pixels. Specifically, we introduce the notion of an action field ${\bm{\mathbf{\Psi}}}=[{\bm{\mathbf{\psi}}}_{1},...,{\bm{\mathbf{\psi}}}_{M}]^{T}$: an $[M\times 2]$ matrix (with $M$ the number of pixels in an image or video frame), such that the $i$-th row in this matrix (${\bm{\mathbf{\psi}}}_{i}$) specifies the (x,y,z)-acceleration applied at pixel $i$. The action on the $k$-th object is then given by:
@@
Refinement network (actions)
Type Size/#Chan. Act. func. Comment Linear 4 LSTM 32 tanh Inputs 10
@@ 6.1.3 Preference network
Encoder
Type Size/#Chan. Act. func. Comment Linear 64 ELU Linear 64 ELU Inputs 64
Decoder (preferences)
Type Size/#Chan. Act. func. Comment Linear 64 ELU Linear 64 ELU Inputs 128
@@ 6.2 Inference procedure
From each belief $q({{\bm{\mathbf{s}}}^{\dagger}}^{(k)})$, we then (as in [10]) sample a state vector (using the reparameterization trick) and run this through the decoder to obtain, for each object, a predicted sub-image ${\bm{\mathbf{o}}}^{(k)}$ and segmentation logits ${\bm{\mathbf{\hat{m}}}}^{(k)}$. These decoder outputs are then fed, separately for each object, into the refinement subnetwork, along with additional inputs. Image-sized inputs are fed into the bottom convolutional layer, while vector-sized inputs are appended to the output of the final convolutional layer and processed by the final LSTM and Linear layers. We use the same selection of auxiliary inputs as [10], except that we omit the leave-one-out likelihoods.
@@ 6.3 Training procedure
The above network architecture was trained on pre-generated experience with the active dSprites environment. The main training set, used to train the world model, comprised 50,000 videos of 4 frames each. An additional validation set of 10,000 videos was sampled identically and independently to the training set. Each video was generated as follows. First, an instance of the active dSprites environment was randomly initialized, with three objects. Object shapes (square, ellipse or heart) were drawn uniformly. Colors were uniformly sampled from 5 evenly spaced values, independently for the R, G and B channels (a uniform grayscale background color was sampled from the same 5 intensity values). Orientations were sampled uniformly from the interval $[0,2\pi]$. Object sizes were sampled uniformly from $[\frac{1}{6}],\frac{1}{3}]$, where $1$ is the size of the image frame. Positions were sampled uniformly from the interval $[0.2,0.8]^{2}$ (where 00 and $1$ are the edges of the frame), while velocities were drawn from a Normal distribution with mean 00 and standard deviation $0.0625$. Video frames were then generated by simulating 4 frames of these objects’ dynamics. Between the 2nd and 3rd frames, a single action (acceleration) was randomly sampled for each object, from a Normal distribution with mean 00 and s.d. $0.0625$, and placed at a random pixel location within the object’s segmentation mask. If an object happened to be completely invisible (e.g. because it was fully occluded by another), then it received no action. An additional action was sampled for a random background pixel, to encourage the model to learn to accurately segment the background into a separate object slot (rather than grouping the background with one of the objects). Training with these 4-frame videos was followed by a brief period of training with longer videos (12 frames, with actions after every second frame) that were otherwise identically sampled.
@@
Additional sets of 50,000 training videos (one for each task; as well as accompanying sets of 10,000 identically and independently sampled validation videos) were used to train OBR’s preference network. Each of these videos was 8 frames long and generated as follows. First, an instance of the active dSPrites environment was randomly initialized as before (but with a different procedure for sampling shapes – see below). Then, 6 frames of this environment were simulated with random actions inserted before the 3rd and 5th frames. To prevent objects from leaving the frame (which is likely in longer videos and becomes problematic in this setting), we additionally placed actions as needed to prevent this from happening (specifically, if an object was on course to move to a coordinate outside the range $[0.05,0.95]^{2}$ in the next frame, an action was generated that resulted in the object appearing to ”bounce” against an invisible wall instead). Before the 7th and 8th frames, actions were generated that brought the objects to their target positions and velocities within the context of a task (see below). The first of these actions brought the objects to their target positions, while the second decelerated them to 0 velocity. Thus, in the final frame (and only then), the objects reached their target configuration. Each frame was encoded (using the normal inference procedure) as a set of latent beliefs in OBR’s latent space, and the preference network was then tasked to map the representation of the first 7 ”seed frames” to that of the final target frame, by minimizing the following KL-divergence loss:
@@
Training was performed using the ADAM optimizer [52] with default parameters and an initial learning rate of $3\times 10^{-4}$. This learning rate was reduced automatically by a factor 3 whenever the validation loss had not decreased in the last 10 training epochs, down to a minimum learning rate of $3\times 10^{-5}$. Training was performed with a batch size of 64 (16 $\times$ 4 GPUs). OBR’s world model was deemed to have converged after 241 epochs, which required approximately 24 hours to train on 4 Nvidia A100 GPUs. Preference networks for the IfHeart and HeartXORSquare tasks were trained for 90 and 120 epochs, respectively (18 and 24 hours).
@@ 6.3.1 ELBO loss
where $\mathcal{H}(\bullet)$ denotes entropy. Similar to previous work (e.g. [10]) we up-weight the reconstruction accuracy term in this loss by a factor $\beta$. We train the network to minimize not just the loss at the end of the inference iterations through the network, but a composite loss that also includes the loss after earlier iterations. Let $\mathcal{L}_{\beta}^{(n)}$ be the loss after $n$ inference iterations, then the composite loss is given by:
@@ 6.3.2 Hyperparameters
OBR includes a total of 4 hyperparameters: (1) the loss-reweighting coefficient $\beta$ (see above); (2) the variance of the pixels around their predicted values, $\sigma_{o}^{2}$; (3) the variance of the noise in the latent space dynamics, $\sigma_{s}^{2}$; and (4) the variance of the noise in the object actions, $\sigma_{\psi}^{2}$. The results described in the current work were achieved with the following settings:
Param. Value $\beta$ 5.0 $\sigma_{o}$ 0.3 $\sigma_{s}$ 0.1 $\sigma_{\psi}$ 0.3
@@
where we slightly abuse notation in the sampling of the pixel assignments, as a vector is sampled from a distribution over a categorical variable. The reason this results in a vector is because this sampling step uses the Gumbel-Softmax trick [53], which is a differentiable method for sampling categorical variables as ”approximately one-hot” vectors. Thus, for every pixel $i$, we sample a vector ${\bm{\mathbf{\hat{m}}}}^{*(i)}_{j}$, such that the $k$-th entry of this vector, $\hat{m}^{*(i)}_{jk}$, denotes the ”soft-binary” condition of whether pixel $i$ belongs to object $k$. In practice, we use $N_{\text{samples}}=1$, based on the intuition that this will still yield a good approximation over many training instances, and that we rely on the refinement network to learn to infer good beliefs. The Gumbel-Softmax sampling method depends on a temperature $\tau$, which we gradually reduce across training epochs, so that the samples gradually better approximate the ideal one-hot vectors.

### v3

@@ 5.2 Grounded protosymbols and preferences in the latent space
A key feature of OBR is that it learns grounded representations and manipulates them in the latent manifold, as well as generating the agent preferences (goals) in the object-centric latent representation. Agents’ internal goals are tied to the world and the actions that it can perform in it. Hence, bringing benefits but also restrictions as it forces the system to perform embodied reasoning.
Learned protosymbolic rules depend on the potential learnable object features and how the agent can act on them—See latent space traversals in Appendix 8. For instance, the agent cannot learn to change the color of an object if all objects are in grey-scale or there is not a policy that can cause the desired effect. As a consequence, conditional behaviours that cannot be interpreted in the objects features space cannot be properly resolved in execution. On the good side, goal-conditioned behaviours are always grounded and therefore the gap between reasoning and control vanishes.
@@ 5.3 Shortcomings and future work
Following the brain inspiration, we employed an iterative inference procedure to endorse adaptation. Modules ablation results show that OBR’s perception backbone (IODINE-dyn) displays better morphology discrimination and adaptation to changes in shape and colors than other state-of-the-art slot attention models. Anyhow, object-centric architectures are still too constrained to the discrimination accuracy of the perception module. While variational iterative inference offers more powerful per-scene reasoning and can refine ambiguous decompositions, amortized approaches may offer faster inference at the cost of reduced shape fidelity, suggesting that future work may benefit from hybrid strategies that balance accuracy and efficiency. Large models could help with the scaling to naturalistic images, but objects manipulation in the latent representation (e.g., JEPA or our OBR approaches) is critical to provide the needed abstraction against noise in the perceptual segmentation. Second, dynamics learning in OBR assumes that actions in the latent representation are linear plus noise. While this could fit a hierarchical interpretation of cognition, it may prevent complex non-linear manipulation dynamics. Ablation results shows that the dynamics learning in OBR can be improved. While OBR can recover from collisions, it cannot plan ahead with them. Third, the achieved agent behaviours are much less complex than of current SOTA LLMs/VLM approaches, such as RT architectures [gu2023rt]. OBR provides, however, goal-conditioned behavioural reasoning, which is a major challenge in the field [sundaresan2024rt]. Our comparison reveals that monolithic architectures like V-JEPA 2-AC  [assran2025v] struggle with conditional rules even when trained with oracle demonstrations, while OBR’s modular design achieves better performance using only self-generated demonstrations. This difference stems from architectural properties. OBR’s preference network proposes context-based desired latent goals, enabling compositional generalization across rules with the same trained mechanism. In contrast, monolithic representations lack interpretable structure in the latent space, limiting their capacity to flexibly recombine learned components for novel behavioral requirements. In any case, increased expressivity of the reasoning should be investigated, for instance, by using OBR as the interface between symbolic (e.g., graph-based, language) and subsymbolic representations. Finally, and very relevant to operationalize this approach for robotics applications, the agent needs to include the body restrictions of performing an action in the world [lanillos2021neuroscience, taniguchi2023world]. This can be done by changing the action field mapping by a robotic arm controller.
@@ Acknowledgments
This work has been funded by the SPIKEFERENCE project, co-funded by the Human Brain Project (HBP) Specific Grant Agreement 3 (ID: 945539). AL was funded by the DeepSelf project, Deutsche Forschungsgemeinschaft (DFG), NUM: 467045002.
@@ 6.1.1 Perceptual inference module
This module used a latent dimension of 16. Note that, in the output of the refinement network, this number is doubled once as each latent belief is encoded by a mean and variance, and then doubled again as we represent (and infer) both the states and their first-order derivatives. In the decoder, the latent dimension is doubled only once, as the state derivatives do not enter into the reconstruction of a video frame. As in [Greff2019a_Iodine], we use a spatial broadcast decoder, meaning that the latent beliefs are copied along a spatial grid with the same dimensions as a video frame, and each latent vector is concatenated with the $(x,y)$ coordinate of its grid location, before passing through a stack of transposed convolution layers. Decoder and refinement network architectures are summarized in the tables below. The refinement network takes in 16 image-sized inputs, which are identical to those used in [Greff2019a_Iodine], except that we omit the leave-one-out likelihoods. Vector-sized inputs join the network after the convolutional stage (which processes only the image-sized inputs), and consist of the variational parameters and (stochastic estimates of) their gradients.
Decoder (states) Type Size/#Chan. Act. func. Comment Input ($\bm{\mathbf{\lambda}}$}) 32 Broadcast 34 Appends coordinate channels ConvT $5\times 5$ 32 ELU ConvT $5\times 5$ 32 ELU ConvT $5\times 5$ 32 ELU ConvT $5\times 5$ 32 ELU ConvT $5\times 5$ 4 Sigmoid or Softmax Outputs RGB + mask
Refinement network (states) Type Size/#Chan. Act. func. Comment Linear 64 LSTM 128 tanh Concat ${\bm{\mathbf{\lambda}}},\nabla_{\bm{\mathbf{\lambda}}}\mathcal{L}]$ 256 Appends vector-sized inputs Linear 128 ELU Flatten 800 Conv $5\times 5$ 32 ELU Conv $5\times 5$ 32 ELU Conv $5\times 5$ 32 ELU Inputs 16
@@
To enable action planning across episodes of non-trivial length, it is important that OBR’s perceptual inference and predictive abilities remain stable across longer time windows. In addition, a great benefit of OBR’s slot-based architecture is that object slots can be added or removed at will, without having to learn additional connection weights [van2023object]. However, it is not a given that performance will be robust to such variations. Here, we test the robustness of OBR’s world model, as measured by segmentation and reconstruction accuracy, to both of these dimensions (Fig. 14). Having been trained on 4-frame videos with 3 objects, OBR generalizes well to longer videos with fewer or more objects present in the scene. A slight drop-off in performance towards later frames in the 12-frame testing videos is easily remedied by a very short bout (3 epochs) of additional training with an additional set of longer (also 12-frame) training videos. Generalization to different numbers of objects is characterized by a slight drop in performance as the number increases, but this might also be (partly) attributed to increased complexity of the resulting images.
@@ 6.1.2 Action inference module
A key challenge in multi-object environments such as the evaluated environment, is that an agent’s internal representation of objects has an unknown (and possibly imperfect) correspondence to the objects in the environment. Even if the object features are inferred with perfect accuracy, the order of the objects in the representation is arbitrary. To solve this correspondence problem, we define an action space in which accelerations can be placed at pixel locations within the environment’s image grid, and objects receive the sum of all accelerations that coincide with their visible pixels. Specifically, we introduce the notion of an action field ${\bm{\mathbf{\Psi}}}=[{\bm{\mathbf{\psi}}}_{1},...,{\bm{\mathbf{\psi}}}_{M}]^{T}$: an $M\times 2]$ matrix (with $M$ the number of pixels in an image or video frame), such that the $i$-th row in this matrix (${\bm{\mathbf{\psi}}}_{i}$) specifies the (x,y,z)-acceleration applied at pixel $i$. The action on the $k$-th object is then given by:
@@
Refinement network (actions) Type Size/#Chan. Act. func. Comment Linear 4 LSTM 32 tanh Inputs 10
@@ 6.1.3 Preference network
Encoder Type Size/#Chan. Act. func. Comment Linear 64 ELU Linear 64 ELU Inputs 64
Decoder (preferences) Type Size/#Chan. Act. func. Comment Linear 64 ELU Linear 64 ELU Inputs 128
@@ 6.2 Inference procedure
From each belief $q({{\bm{\mathbf{s}}}^{\dagger}}^{(k)})$, we then (as in [Greff2019a_Iodine]) sample a state vector (using the reparameterization trick) and run this through the decoder to obtain, for each object, a predicted sub-image ${\bm{\mathbf{o}}}^{(k)}$ and segmentation logits ${\bm{\mathbf{\hat{m}}}}^{(k)}$. These decoder outputs are then fed, separately for each object, into the refinement subnetwork, along with additional inputs. Image-sized inputs are fed into the bottom convolutional layer, while vector-sized inputs are appended to the output of the final convolutional layer and processed by the final LSTM and Linear layers. We use the same selection of auxiliary inputs as [Greff2019a_Iodine], except that we omit the leave-one-out likelihoods.
@@ 6.3 Training procedure
The above network architecture was trained on pre-generated experience with the active dSprites environment. The main training set, used to train the world model, comprised 50,000 videos of 4 frames each. An additional validation set of 10,000 videos was sampled identically and independently to the training set. Each video was generated as follows. First, an instance of the active dSprites environment was randomly initialized, with three objects. Object shapes (square, ellipse or heart) were drawn uniformly. Colors were uniformly sampled from 5 evenly spaced values, independently for the R, G and B channels (a uniform grayscale background color was sampled from the same 5 intensity values). Orientations were sampled uniformly from the interval $0,2\pi]$. Object sizes were sampled uniformly from $\frac{1}{6}],\frac{1}{3}]$, where $1$ is the size of the image frame. Positions were sampled uniformly from the interval $0.2,0.8]^{2}$ (where 0 and $1$ are the edges of the frame), while velocities were drawn from a Normal distribution with mean 0 and standard deviation $0.0625$. Video frames were then generated by simulating 4 frames of these objects’ dynamics. Between the 2nd and 3rd frames, a single action (acceleration) was randomly sampled for each object, from a Normal distribution with mean 0 and s.d. $0.0625$, and placed at a random pixel location within the object’s segmentation mask. If an object happened to be completely invisible (e.g. because it was fully occluded by another), then it received no action. An additional action was sampled for a random background pixel, to encourage the model to learn to accurately segment the background into a separate object slot (rather than grouping the background with one of the objects). Training with these 4-frame videos was followed by a brief period of training with longer videos (12 frames, with actions after every second frame) that were otherwise identically sampled.
@@
Additional sets of 50,000 training videos (one for each task; as well as accompanying sets of 10,000 identically and independently sampled validation videos) were used to train OBR’s preference network. Each of these videos was 8 frames long and generated as follows. First, an instance of the active dSPrites environment was randomly initialized as before (but with a different procedure for sampling shapes – see below). Then, 6 frames of this environment were simulated with random actions inserted before the 3rd and 5th frames. To prevent objects from leaving the frame (which is likely in longer videos and becomes problematic in this setting), we additionally placed actions as needed to prevent this from happening (specifically, if an object was on course to move to a coordinate outside the range $0.05,0.95]^{2}$ in the next frame, an action was generated that resulted in the object appearing to ”bounce” against an invisible wall instead). Before the 7th and 8th frames, actions were generated that brought the objects to their target positions and velocities within the context of a task (see below). The first of these actions brought the objects to their target positions, while the second decelerated them to 0 velocity. Thus, in the final frame (and only then), the objects reached their target configuration. Each frame was encoded (using the normal inference procedure) as a set of latent beliefs in OBR’s latent space, and the preference network was then tasked to map the representation of the first 7 ”seed frames” to that of the final target frame, by minimizing the following KL-divergence loss:
@@
Training was performed using the ADAM optimizer [kingma2014adam] with default parameters and an initial learning rate of $3\times 10^{-4}$. This learning rate was reduced automatically by a factor 3 whenever the validation loss had not decreased in the last 10 training epochs, down to a minimum learning rate of $3\times 10^{-5}$. Training was performed with a batch size of 64 (16 $\times$ 4 GPUs). OBR’s world model was deemed to have converged after 241 epochs, which required approximately 24 hours to train on 4 Nvidia A100 GPUs. Preference networks for the IfHeart and HeartXORSquare tasks were trained for 90 and 120 epochs, respectively (18 and 24 hours).
@@ 6.3.1 ELBO loss
where $\mathcal{H}(\bullet)$ denotes entropy. Similar to previous work (e.g. [Greff2019a_Iodine]) we up-weight the reconstruction accuracy term in this loss by a factor $\beta$. We train the network to minimize not just the loss at the end of the inference iterations through the network, but a composite loss that also includes the loss after earlier iterations. Let $\mathcal{L}_{\beta}^{(n)}$ be the loss after $n$ inference iterations, then the composite loss is given by:
@@ 6.3.2 Hyperparameters
OBR includes a total of 4 hyperparameters: (1) the loss-reweighting coefficient $\beta$ (see above); (2) the variance of the pixels around their predicted values, $\sigma_{o}^{2}$; (3) the variance of the noise in the latent space dynamics, $\sigma_{s}^{2}$; and (4) the variance of the noise in the object actions, $\sigma_{\psi}^{2}$. The results described in the current work were achieved with the following settings: Param. Value $\beta$ 5.0 $\sigma_{o}$ 0.3 $\sigma_{s}$ 0.1 $\sigma_{\psi}$ 0.3
@@
where we slightly abuse notation in the sampling of the pixel assignments, as a vector is sampled from a distribution over a categorical variable. The reason this results in a vector is because this sampling step uses the Gumbel-Softmax trick [Jang2016gumbel], which is a differentiable method for sampling categorical variables as ”approximately one-hot” vectors. Thus, for every pixel $i$, we sample a vector ${\bm{\mathbf{\hat{m}}}}^{*(i)}_{j}$, such that the $k$-th entry of this vector, $\hat{m}^{*(i)}_{jk}$, denotes the ”soft-binary” condition of whether pixel $i$ belongs to object $k$. In practice, we use $N_{\text{samples}}=1$, based on the intuition that this will still yield a good approximation over many training instances, and that we rely on the refinement network to learn to infer good beliefs. The Gumbel-Softmax sampling method depends on a temperature $\tau$, which we gradually reduce across training epochs, so that the samples gradually better approximate the ideal one-hot vectors.